Consider a Keynesian model:

Full employment output = R100 million

Tax rate = 0,25

Investment = R40 million

Autonomous consumption = R30 million

Marginal propensity to consume = 0,8


The value of the multiplier is …

[1] 2

[2] 1.67

[3] 2.5

[4] 4


The equilibrium level of income is …

[1] R70 million.

[2] R175 million.

[3] R280 million.

[4] R140 million.


To bring about full employment, government spending should be …

[1] -R30 million.

[2] -R72 million.

[3] R30 million.

[4] R75 million.


1
Expert's answer
2021-06-29T10:45:06-0400

Solution:

a.). The value of the multiplier is [4.] 4.

Multiplier = "\frac{1}{(1 - MPC)}"

MPC = 0.8

Multiplier = "\frac{1}{(1 - 0.8)} = \frac{1}{0.2} = 5"

The value of the multiplier = 5

 

b.). The equilibrium level of income is [2.] R175 million.

At equilibrium: AD = AS

Y = C + I + G

C = 30 + 0.8YD = 30 + 0.8(Y – T) = 30 + 0.8(Y – 0.25Y)

C = 30 + 0.8(Y – 0.25Y)

I = 40

Y = 30 + 0.8(Y – 0.25Y) + 40

Y = 30 + 0.6Y + 40

Y – 0.6Y = 30 + 40

0.4Y = 70

Y = 175

The equilibrium level of income is R175 million.

 

c.). To bring about full employment, government spending should be [1] -R30 million.

GDP needs to be decreased by (175 – 100 = 75 million) to bring about full employment.

Plug the full employment level of 100 million same as Y and solve for government spending.

Y = C + I + G

100 = 30 + 0.8(Y – 0.25Y) + 40 + G

100 = 30 + 0.8(100 – 0.25(100) + 40 + G

100 = 30 + 60 + 40 + G

100 = 130 + G

100 – 130 = G

G = -30

To bring about full employment, government spending should be -R30 million.